Find the curvature, the radius and the center of curvature at a point.


r=1+ cos theta ,theta=Ï€/2

1
Expert's answer
2022-01-18T11:37:49-0500

"r=f(\theta)=1+cos\theta \Rightarrow f(\frac{\pi}{2})=1\\
f'(\theta)=-sin\theta\Rightarrow f'(\frac{\pi}{2})=-1\\
f''(\theta)=-cos\theta \Rightarrow f''(\frac{\pi}{2})=0\\
\therefore \text{Radius of curvature}\\= \dfrac{(f(\frac{\pi}{2})^2+f'(\frac{\pi}{2})^2)^\frac{3}{2}}
{f(\frac{\pi}{2})^2+2f'(\frac{\pi}{2})^2-f(\frac{\pi}{2})f''(\frac{\pi}{2})}\\
=\dfrac{2^\frac{3}{2}}{3}=\dfrac{2\sqrt 2}{3}\\
\text{Now, curvature $=$}\dfrac{1}{\text{Radius of Curvature}}=\dfrac{3}{2\sqrt 2}"



"\text{And, Center of curvature = }((y_0+R),(x_0+f'(x_0))R)\\
\texttt{[Where $x_0=\frac{\pi}{2}, y_0=f(\frac{\pi}{2})=1$]}\\
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=
(1+ \frac{2\sqrt 2}{3}, (\frac{\pi}{2}-1)\frac{2\sqrt 2}{3})\\"